# Keeping Up with the Models: Online Deployment and Routing of LLMs at Scale

### v1

@@ B.5 Total Regret Bound
$\mathcal{R}(T)\leq\mathcal{O}\left(\sqrt{M_{\max}KT\log(NT/\delta)}+\frac{NT}{M_{\max}K}\right)$
This composite regret bound highlights different facets of the online learning problem. The first two terms capture the cost of learning model parameters (both for routing and deployment) and adapting the deployment strategy within a limited number of concurrently active models ($M_{\max}$) across $K$ stages. The third term, $\mathcal{O}(NT/(M_{\max}K))$, explicitly incorporates the penalty associated with the breadth of the model universe ($N$) relative to deployment capacity ($M_{\max}$) and the granularity of adaptation opportunities ($K$ stages). This term becomes significant if many models need to be evaluated through a narrow deployment bottleneck over a limited number of stages.
Appendix C Lower Bound for Online LLM Routing
@@
For each batch $j\in\{1,\dots,N_{B}\}$, we consider a set of $M$ models available to the algorithm $\mathcal{A}$. These models are indexed $i\in\{1,\dots,M\}$ specifically for the current batch $j$. The performance characteristics of these $M$ models are defined as follows: one model is designated as the “strong” model, and the remaining $M-1$ models are “weak”. Let $s_{j}\in\{1,\dots,M\}$ be the index of this strong model in batch $j$, chosen uniformly at random by the adversary and unknown to the algorithm. The reward $r_{t}$ obtained from selecting a model at query $t\in\mathcal{B}_{j}$ is drawn from a Bernoulli distribution:
@@
Fix a batch $j\in\{1,\dots,N_{B}\}$, and let $s_{j}^{*}$ denote the true index of the strong model, unknown to algorithm $\mathcal{A}$. To identify $m_{s_{j}^{*}}$ (mean $\mu_{H}^{(j)}$) from the $M-1$ weak models (mean $\mu_{L}^{(j)}$, gap $\epsilon$) with a constant probability $p_{succ}<1$ (e.g., $p_{succ}=3/4$), any algorithm requires a query complexity of $\Omega(M\epsilon^{-2})$ [2, 30]. This implies there is a universal constant $c_{S}>0$ such that at least $c_{S}M\epsilon^{-2}$ queries are necessary to achieve success probability $p_{succ}$.
@@
    Value of $\Delta$: $\epsilon^{2}=\left(\frac{c_{S}M}{8T}\right)^{2/3}$. $\Delta=(c_{S}/2)M\epsilon^{-2}=(c_{S}/2)M\left(\frac{c_{S}M}{8T}\right)^{-2/3}=(c_{S}/2)M\frac{(8T)^{2/3}}{(c_{S}M)^{2/3}}$ $=(c_{S}/2)M\frac{8^{2/3}T^{2/3}}{c_{S}^{2/3}M^{2/3}}=(c_{S}/2)M\frac{4T^{2/3}}{c_{S}^{2/3}M^{2/3}}=\frac{c_{S}}{2}\frac{4M^{1/3}T^{2/3}}{c_{S}^{2/3}}=2c_{S}^{1/3}M^{1/3}T^{2/3}$. Let $c_{\Delta}=2c_{S}^{1/3}$. So $\Delta=c_{\Delta}M^{1/3}T^{2/3}$. We need $1\leq\Delta\leq T$ for $\Delta$ to be a valid batch length. $\Delta\leq T\implies c_{\Delta}M^{1/3}T^{2/3}\leq T\implies M^{1/3}\leq T^{1/3}/c_{\Delta}\implies T\geq(c_{\Delta}M^{1/3})^{3}=c_{\Delta}^{3}M$. This implies $T$ must be sufficiently large relative to $M$. $\Delta\geq 1$ generally holds for large $T$ if $M\geq 1$.

### v2

@@ C.5 Total Regret Bound
$\mathcal{R}(T)\leq\mathcal{O}\left(\sqrt{M_{\max}KT\log(NT/\delta)}+\frac{NT}{M_{\max}K}\right).$
Appendix D Lower Bound for Online LLM Routing
@@
For each batch $j\in\{1,\dots,N_{B}\}$, we consider a set of $M$ models available to the algorithm $\mathcal{A}$. These models are indexed $i\in\{1,\dots,M\}$ specifically for the current batch $j$. The performance characteristics of these $M$ models are defined as follows: one model is designated as the “strong” model, and the remaining M1M-1 models are “weak”. Let $s_{j}\in\{1,\dots,M\}$ be the index of this strong model in batch $j$, chosen uniformly at random by the adversary and unknown to the algorithm. The reward $r_{t}$ obtained from selecting a model at query $t\in\mathcal{B}_{j}$ is drawn from a Bernoulli distribution:
@@
Fix a batch $j\in\{1,\dots,N_{B}\}$, and let $s_{j}^{*}$ denote the true index of the strong model, unknown to algorithm $\mathcal{A}$. To identify $m_{s_{j}^{*}}$ (mean $\mu_{H}^{(j)}$) from the M1M-1 weak models (mean $\mu_{L}^{(j)}$, gap $\epsilon$) with a constant probability $p_{succ}<1$ (e.g., $p_{succ}=3/4$), any algorithm requires a query complexity of $\Omega(M\epsilon^{-2})$ Agarwal et al. (2017); Li et al. (2023). This implies there is a universal constant $c_{S}>0$ such that at least $c_{S}M\epsilon^{-2}$ queries are necessary to achieve success probability $p_{succ}$.
@@
    Value of $\Delta$: $\epsilon^{2}=\left(\frac{c_{S}M}{8T}\right)^{2/3}$. $\Delta=(c_{S}/2)M\epsilon^{-2}=(c_{S}/2)M\left(\frac{c_{S}M}{8T}\right)^{-2/3}=(c_{S}/2)M\frac{(8T)^{2/3}}{(c_{S}M)^{2/3}}$ $(c_{S}/2)M\frac{8^{2/3}T^{2/3}}{c_{S}^{2/3}M^{2/3}}=(c_{S}/2)M\frac{4T^{2/3}}{c_{S}^{2/3}M^{2/3}}=\frac{c_{S}}{2}\frac{4M^{1/3}T^{2/3}}{c_{S}^{2/3}}=2c_{S}^{1/3}M^{1/3}T^{2/3}$. Let $c_{\Delta}=2c_{S}^{1/3}$. So $\Delta=c_{\Delta}M^{1/3}T^{2/3}$. We need $1\leq\Delta\leq T$ for $\Delta$ to be a valid batch length. $\Delta\leq T\implies c_{\Delta}M^{1/3}T^{2/3}\leq T\implies M^{1/3}\leq T^{1/3}/c_{\Delta}\implies T\geq(c_{\Delta}M^{1/3})^{3}=c_{\Delta}^{3}M$. This implies $T$ must be sufficiently large relative to $M$. $\Delta\geq 1$ generally holds for large $T$ if $M\geq 1$.
@@
Appendix E Experimental Details
This appendix provides additional details on the experimental setup, benchmarks, and results presented in Section 5.
E.1 Real-World Benchmark: RouterBench
Dataset Details. Our primary evaluation is based on the RouterBench dataset (Hu et al., 2024), a comprehensive benchmark with 36,497 queries sampled from eight diverse NLP datasets. These datasets cover both Chinese and English and span a broad spectrum of tasks: commonsense reasoning (HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC Challenge (Clark et al., 2018)), knowledge-based understanding (MMLU (Hendrycks et al., 2021)), open-domain dialogue (MT-Bench (Zheng et al., 2023)), mathematical reasoning (GSM8K (Cobbe et al., 2021)), code generation (MBPP (Austin et al., 2021)), and retrieval-augmented generation (RAG).
Candidate LLMs and Metrics. For each query, RouterBench provides pre-computed responses from the 11 LLMs listed in Table 3. The table shows the models ordered by their release date, along with their average performance score and average cost per query across the entire dataset.
Model Avg Performance Avg Cost
claude-instant-v1 0.5900 \$0.001236
claude-v1 0.6480 \$0.005870
claude-v2 0.5116 \$0.006153
meta/llama-2-70b-chat 0.6059 \$0.001337
WizardLM/WizardLM-13B-V1.2 0.5392 \$0.000142
meta/code-llama-instruct-34b-chat 0.5040 \$0.000550
mistralai/mistral-7b-chat 0.4999 \$0.000139
gpt-3.5-turbo-1106 0.6867 \$0.000709
gpt-4-1106-preview 0.8048 \$0.007943
zero-one-ai/Yi-34B-Chat 0.7153 \$0.000558
mistralai/mixtral-8x7b-chat 0.6504 \$0.000414
E.2 Additional Simulation Results
To validate our framework’s applicability to the rapidly evolving frontier of SOTA models, we constructed a synthetic benchmark using 15 recent, high-performance LLMs.
Candidate LLMs and Metrics. We consider a set of 15 LLMs, summarized in Table 4, ordered by their official release dates. For each model, we report both performance metrics  Chiang et al. (2024) and associated costs. No single metric can capture the full complexity of LLM performance. We chose Elo / area scores as they are community standards (e.g., Chatbot Arena), ensuring transparency and reproducibility. Importantly, our framework is metric-agnostic. The performance signal can be replaced by any other quantifiable performance measure, such as task-specific accuracy, user satisfaction scores, or a composite utility function combining multiple objectives.
Note that the naming of AI models can be highly nuanced, and multiple versions may exist under a similar label. For example, for Gemini 2.5 Pro, we used data corresponding to the Gemini-2.5-Pro-Preview-05-06 version. Additionally, performance scores may fluctuate over time due to model updates, shifts in the user voting population, or changes in evaluation benchmarks. Similarly, costs may vary across time or across versions. The data reported here corresponds to the specific versions and conditions used in our experiments.
Simulation Setup. These 15 models are introduced sequentially, with the first 5 available at time $t=1$ and one additional model becoming available every 1,000 rounds until all are accessible. Each model’s performance is normalized to lie within the $[0,1]$ interval. For each model $m$, we set its budget weight parameter $\alpha_{m}=0.4$, except for Yi-Lightning, for which we set $\alpha_{m}=1$. This is because Yi-Lightning is the cheapest model initially available to the algorithm, ensuring the feasibility of the mixed-integer optimization problem in every round. At each round, if the algorithm selects a model, it receives a noisy performance and cost observation, where the noise is sampled from a Gaussian distribution centered around the true value. In any real-world application, LLM performance metrics are inherently bounded. Our simulation implicitly respects the $[0,1]$ range by clipping noisy rewards to this interval, with the noise variance chosen such that the probability of generating values outside the range is negligible.
(a) $M_{\max}$ = 3.
(b) $M_{\max}$ = 5.
(c) $M_{\max}$ = 10.
(d) Update interval = 250 rounds.
(e) Update interval = 1000 rounds.
(f) $b$ = 5.
Figure 6: Regret with varying parameters.
Sensitivity Analysis. Figure 6 presents the sensitivity analysis for StageRoute on the synthetic benchmark. The results are consistent with our findings on RouterBench, demonstrating the robustness of our algorithm across different model suites and settings. StageRoute consistently achieves low regret, adapting effectively to the concurrency cap, update interval, and budget.
Table 4: LLM comparison by release date: performance (Arena Score/Elo Chiang et al. (2024)) and output cost (per 1M tokens).
Model Performance Cost
GPT-4o 1336 \$10.00
Gemini-1.5-pro-002 1302 \$5.00
Yi-Lightning 1287 \$0.14
o1-mini 1304 \$0.60
Llama 3.3 70B Instruct 1257 \$0.40
DeepSeek-R1 1358 \$0.55
Gemini 2.0 Flash 1380 \$1.50
Claude 3.7 Sonnet 1300 \$15.00
Hunyuan-turbos 1296 \$0.28
Deepseek-v3 1373 \$0.28
Llama-4-Maveric 1417 \$0.82
GPT-4.1 1366 \$8.00
Grok-3-preview 1403 \$15.00
o3 by OpenAI 1413 \$40.00
Gemini-2.5-Pro 1446 \$10.00